calculate the frequency of the fundamental note produced by a string 1 m long and weighing 2 g kept stretched by a load 400 kg

Grade:12th pass

3 Answers

Elton
7 years ago
direct substitution answer.. :) f=?, l=1m, T(tension)=400N, m(=mass/length)=2*10-3
f=1/2l({\sqrt{T/m}})...substitute..u get the answer.. :)
Gosula Madhukar
7 years ago
the fundamental frequency is
=[(200000)^1/2]/2
=223.6 hz
I think iam correct.
if u have any further doubts please post them.
thank u.
 
Physics
3 years ago
Tension is mass multiplied by acceleration due to to gravity. So using formula f=1/2L (T/m) ^1/2. Where m is mass per unit length
the answer will be around 700Hz